# Task Prompt Vectors: Effective Initialization through Multi-Task Soft-Prompt Transfer

### v1

@@ Abstract
Prompt tuning is a modular and efficient solution for training large language models (LLMs). One of its main advantages is task modularity, making it suitable for multi-task problems. However, current soft-prompt-based methods often sacrifice multi-task modularity, requiring the training process to be fully or partially repeated for each newly added task. While recent work on task vectors applied arithmetic operations on full model weights to achieve the desired multi-task performance, a similar approach for soft-prompts is still missing. To this end, we introduce Task Prompt Vectors, created by element-wise difference between weights of tuned soft-prompts and their random initialization. Experimental results on 12 NLU datasets show that task prompt vectors can be used in low-resource settings to effectively initialize prompt tuning on similar tasks. In addition, we show that task prompt vectors are independent of the random initialization of prompt tuning. This allows prompt arithmetics with the pre-trained vectors from different tasks. In this way, by arithmetic addition of task prompt vectors from multiple tasks, we are able to outperform a state-of-the-art baseline in some cases.
@@ 1 Introduction
Figure 1: An illustration of task prompt vector and the combination via addition that we include in our work. (a) A task prompt vector is created by subtracting the soft-prompt initialization weights $\theta_{P_{pre}}$ from the soft-prompt weights after prompt tuning $\theta^{t}_{P_{ft}}$ (Section 3, eq. 2). (b) A combination via the addition of two task prompt $\tau_{P_{a}}$ and $\tau_{P_{b}}$ resulting in $\tau_{P_{new}}$ (Section 3, eq. 4). (c) Different task prompt vectors pointing into the same sub-space in the embedding space of PLM (Section 3.1). The circles represent different random initializations.
@@
To address the problem of the increasing number of parameters, Parameter-Efficient Fine-Tuning (PEFT) methods (Lester et al., 2021; Houlsby et al., 2019; Hu et al., 2022) were introduced, capable of solving multiple problems even with small amounts of labeled data while training only a fraction of the model parameters (e.g., for RoBERTa base (Liu et al., 2019), prompt tuning Lester et al. (2021) is training only 0.5% parameters, and LoRA (Hu et al., 2022) is training only 0.7% of parameters (Xu et al., 2023)). The key concept that makes such methods effective is their task modularity – single modules can be trained for diverse tasks or languages and then just be swapped out inside of the same model.
@@
Some of the recent PEFT (Xu et al., 2023; Lester et al., 2021; Asai et al., 2022) methods focus on fine-tuning soft-prompts. Soft-prompts are trainable (parametrized) weights that are prepended to the input embeddings while training the model. Prompt tuning is one such modular and efficient solution for soft-prompt-based tuning of large language models (LLMs), with an advantage in task modularity, making it suitable for multi-task problems. Prompt tuning scales linearly only with the internal size of model embeddings not with the total number of parameters, compared to other PEFT methods (the number of parameters is consistent across different model sizes), making it exceptionally efficient, while matching the performance of fully fine-tuned models.
@@
Current soft-prompt-based methods often lack multi-task modularity, requiring the training process to be fully or partially repeated for each newly added task (Vu et al., 2022; Wang et al., 2023). Other methods while keeping their relatively high modularity usually lack robustness and their performance depends on the quality and the number of pre-trained soft-prompts (Asai et al., 2022). Moreover, creating a soft-prompt for multiple tasks may often reduce the overall multi-task performance and require further fine-tuning. Building upon the findings from task vector arithmetics Ilharco et al. (2022), we utilize the efficiency and modularity of prompt tuning Lester et al. (2021) to create Task Prompt Vectors. We thoroughly investigate the properties of task prompt vectors and demonstrate their functionality in combining pairs of task prompt vectors while evaluating their in-distribution performance and out-of-distribution performance in full and limited data scenarios.
@@
Our main contributions and findings are  [^1] To support the replicability of our work, we provide a repository where we store all of our implementation and results: https://github.com/Wicwik/task-prompt-vectors:
@@ 2 Related Work
To the best of our knowledge, there is no research on task vectors in the context of soft-prompt-based fine-tuning. In this work, we address this drawback by combining the existing knowledge on prompt tuning and task vectors.
3 Method
3.1 Background
In this section, we describe the necessary background to Prompt Tuning and Task Vectors that we will build on in Section 3.2.
Prompt tuning.
Prompt tuning (Lester et al., 2021) is a widely used parameter-efficient fine-tuning (PEFT) method for fine-tuning a pre-trained language model (PLM) on a variety of downstream tasks. It is conceptually based on the prompting technique, i.e., the interaction with an LLM via text. Instead of using tokenized words, soft-prompts are continuous representations that are passed to an LLM. Prompt tuning, as introduced in Lester et al. (2021), casts all tasks as text generation, modeling a probability $Pr(Y|X)$, where $X$ is a sequence of input tokens and $Y$ is a sequence of output tokens representing the class label. The classification $Pr_{\theta}(Y|X)$ is then parametrized by the model weights $\theta$. Prompting adds extra information to the classification process by prepending a series of tokens (prompt) $P$ to the input $X$, such that the model maximizes the probability of getting current $Y$ in $Pr_{\theta}(Y|[P;X])$, while keeping the parameters $\theta$ frozen. Prompt tuning adds another parameter $\theta_{P}$ to the equation, which parametrizes the prompt. During the training, only $\theta_{P}$ is updated as the following function is optimized:
$\mathcal{L}_{PT}=-\sum_{i}logPr_{\theta,\theta_{P}}(Y_{i}|[P;X_{i}])$ (1)
Task vectors and task arithmetics.
As a method of editing model weights without training, task vectors (Ilharco et al., 2022) assume the existence of pre-trained weights $\theta_{pre}\in\mathbb{R}^{d}$. Weights of the same model after fine-tuning on a downstream task $t$ are represented as $\theta^{t}_{ft}\in\mathbb{R}^{d}$. Task vector $\tau_{t}\in\mathbb{R}^{d}$ is given by the element-wise difference:
$\tau_{t}=\theta^{t}_{ft}-\theta_{pre}$ (2)
Task vectors can be then applied to any model weights $\theta$ of the same dimensionality (architecture) by element-wise addition with a rescaling term:
$\theta_{new}=\theta+\lambda\tau_{t}$ (3)
The rescaling term $\lambda$ is a real number with respect to $0<\lambda\leq 1$ and when $\lambda=1$ then $\theta_{new}=\theta^{t}_{ft}$.
The task vector representation in the weight space of the model has the same properties as standard vectors, therefore it is possible to include them in arithmetic expressions like addition, negation, or combinations via the addition of two or more vectors together:
$\tau_{new}=\sum_{i}\tau_{i}$ (4)
3.2 Task Prompt Vectors
In this work, we proceed from the concepts originating in the parameter-efficient prompt tuning and task vectors. Let $T_{1},...,T_{t}$ be a set of source tasks and $\theta_{P_{1}},...,\theta_{P_{i}}$be a set of random soft-prompt weights initializations. Intuitively, the random soft-prompt weights initializations are random points in the embedding space of the PLM. We then move each of these points (via prompt tuning) into a task sub-space where the optimization function from the equation 1 returns the (sufficiently) minimal value and we repeat this for each task $t\in T$. These points are further denoted as task prompts – soft-prompts fine-tuned by prompt tuning to a set of downstream tasks. The straight trajectory from the initial point to the task prompt is our task prompt vector (see Figure 1 part a)).
@@ 3.2 Task Prompt Vectors
Let $\theta_{P_{pre}}\in\mathbb{R}^{d}$ be the weights of the soft-prompt randomly initialized from the embedding vocabulary of a PLM, and $\theta^{t}_{P_{ft}}\in\mathbb{R}^{d}$ be the weights of the soft-prompt $P$ fine-tuned on a specific task $t$, using the standard prompt tuning formula from equation 1. We formulate the task prompt vector $\tau_{P_{t}}$ for soft-prompt $P$ and task as an element-wise difference, according to the equation 2:
$\tau_{P_{t}}=\theta^{t}_{P_{ft}}-\theta_{P_{pre}}$ (5)
Similar to task vectors, applying a task prompt vector to the soft-prompt weights of the same size would follow the equation 3:
$\theta_{P_{new}}=\theta_{P}+\lambda\theta^{t}_{P_{ft}}$ (6)
Where the rescaling term $\lambda$ is a number from the same interval $0<\lambda\leq 1$ and when $\lambda=1$, then $\theta_{P_{new}}=\theta^{t}_{P_{ft}}=\theta_{P}$.
Task prompt vector arithmetic.
The task prompt vectors for different tasks can be combined by a simple vector addition arithmetic, combining knowledge from different tasks. When we experiment specifically with combinations of two different task prompt vectors, we can rewrite equation 4 into the additions of two task prompt vectors (see Figure 1 part b)):
$\tau_{P_{new}}=\tau_{P_{a}}+\tau_{P_{b}}$ (7)
@@ 4.1 Experimental Setup and Implementation Details
We investigate the properties of task prompt vectors using the representative foundation T5-base (Raffel et al., 2020) model. Our investigation covers 3 types of classification problems and 12 corresponding datasets, namely natural language inference (MNLI (Williams et al., 2018), QNLI (Wang et al., 2018), SciTail (Khot et al., 2018), SNLI (Bowman et al., 2015)), topic classification (DBPedia (Auer et al., 2007), TREC Coarse (Li and Roth, 2002; Hovy et al., 2001), AG News, Yahoo Answers (Zhang et al., 2015)), and sentiment classification (SST2 (Socher et al., 2013), Yelp Polarity, SST5, IMDB (Maas et al., 2011)). Half of the datasets are used as source ones (to obtain pre-trained task prompt vectors), while the second half is purposefully left as target datasets, on which the pre-trained vectors are applied to. In the scope of our work, we refer to a single dataset as a task.
For all experiment results, we report F1 macro, if not specified otherwise. The cosine similarity between vectors (task prompts or task prompt vectors) is measured using the flattened weights of each vector (which has a size of $100\times 768$ parameters, resulting in a 76800-dimensional vector). Since we are utilizing the T5-base for conditional generation, we are computing exact match instead of accuracy for classification. Because we are generating labels also for classification tasks, the exact match is equivalent to accuracy in the sequence classification task. We average our zero- and few-shot results across 3 different runs (i.e., different random initializations of soft-prompts) for ATTEMPT and multi-task SPoT baselines (mostly to save more computational resources) and across 10 different runs for all other experiments. To determine the statistical significance of our results we perform a two-sample Student’s t-test (Student, 1908) with Bonferroni correction Dunn (1959) between the best result and the second best result. If the population sizes differ (e.g. 10 and 3 runs) we use Welch’s t-test (Welch, 1947).
We set soft-prompt length to 100 tokens, learning rate to 0.3, and lower the weight decay of the AdamW optimizer (Loshchilov and Hutter, 2019) to $1\times 10^{-5}$. We train all models on all data sets for 10 epochs, except for TREC, where we train for 50 epochs due to the tendency of models to underfit here. For the few-shot experiments (simulating limited labeled data scenarios), we randomly sub-sample from the data for the respective number of shots while keeping the class distribution. We train for 1000 update steps while keeping a batch size of 2 for 5, 10, 25 shots, 8 for 50, 100, 250 shots, and 16 for 500, 750, 1000 shots. In our work, we consider shot and sample to be equivalent (i.e., for a 5-shot setting we choose 5 samples overall, and not 5 samples per class). A more detailed description of our experimental setup and hyperparameters description can be found in Appendix A.
@@ 4.2 Investigating Task Prompt Vectors Properties
dataset QNLI MNLI TREC Coarse DBpedia SST2 Yelp avg
Prompt tuning $93.3_{0}$ $85.4_{0.1}$ $95.5_{1.7}$ $99.1_{0}$ $93.8_{0.3}$ $97.2_{0}$ $93.8_{0.08}$
Random init ${93.2_{0.1}}^{*}$ ${85.3_{0.2}}$ ${26.5_{18.2}}^{*}$ ${99_{0.1}}^{*}$ ${93.2_{0.6}}$ ${97.1_{0.1}}^{*}$ ${82.4_{3.2}}^{*}$
Table 1: Comparison of exact match results across 10 random soft-prompt initializations. The subscript represents the standard deviation from the average. The first row (Prompt tuning) represents the results of prompt tuning averaged across 10 random initializations. The second row (Random init) represents the results of applying a task prompt vector created from different random initializations to a specific random initialization for all 10 random initializations (covering all 100 random initialization combinations). The * in the superscript represents the results where two-sample Student’s t-test (Student, 1908) confirmed the statistical significance.
RQ1: How universally can we apply task prompt vectors to a) different prompt initialization and b) different tasks?
@@
To evaluate these properties, we train a set of soft-prompts on specified source tasks for inference classification (MNLI, QNLI), topic classification (DBPedia, TREC Coarse), and sentiment classification (SST2, Yelp Polarity), resulting in a set of six soft-prompts that were trained from a single random initialization. We sample 10 random initializations from which we create the task prompt vectors as described in Section 3. We aggregate by averaging our results across random initializations in Table 1 and Figures 2, 3. We start with the evaluation of whether the task prompt vectors are independent of the random initialization and continue with the experiments to confirm whether the trained task prompts from prompt tuning end up in the same task sub-space of the PLM embedding space. This helps us determine whether the task prompt vectors point in the same space, similar to Figure 1 part c).
@@
We conduct experiments to evaluate the performance of applying task prompt vectors to different random initializations. For each task and each random initialization, we apply the task prompt vector (according to the equation 3) to all of the other random initializations and evaluate performance for each task prompt vector-initialization pair on the test set of the particular dataset. Row 2 in Table 1 shows the aggregated exact match results. The results differ only slightly in most observed tasks, compared to the results of prompt tuning. This indicates that task prompt vectors perform well irrespective of their initialization. The only exception is in the TREC task, where the performance decreases drastically. We suspect that this may be caused by the task being harder for the T5-base model to learn, which also confirms the higher standard deviation from the mean of prompt tuning performance.
@@
Figures 2 and 3 show the comparison of cosine similarities between task prompts and task prompt vectors from different tasks averaged over all random initialization combinations. We can see from the low cosine similarities in both tables, that the task prompts and task prompt vectors do not end up in the same space/direction when initialized from different points in the embedding space. The highest cosine similarities on the diagonal represent the highest cosine similarity that we can get by comparing vectors of the same task only across the different random initializations, which serves as a baseline for comparison with the cross-task cosine similarities. We can see in Table 1 row 1, that the downstream performance of prompt tuning on the source tasks across 10 different random initializations has a low standard deviation from the average. This means that the task prompts after prompt tunings end up in a subspace with sufficient task performance, without necessarily pointing to the same spot in the task subspace. Cosine similarities that we have used to create the aggregated figures can be seen in the Appendix B in Figures 6 and 8.
Figure 2: Comparison of average cosine similarities of task prompts fine-tuned on different tasks. The average is calculated across all combinations of 10 random initializations (i.e., row QNLI column MNLI was calculated as the average of all cosine similarities between MNLI and QNLI task prompts for all random initialization combinations omitting the combinations where cosine similarity is equal to 1). The diagonal represents the cosine similarities across different random initializations of the same tasks and it represents the maximum value of cosine similarity if we keep the tasks frozen and compare only across different random initializations.
Figure 3: Comparison of average cosine similarities of task prompt vectors. The averages are calculated similarly to Figure 2 but with task prompt vectors created from different task prompts. The diagonal represents the cosine similarities across different random initializations of the same tasks and it represents the maximum value of cosine similarity if we keep the tasks frozen and compare only across different random initializations.
Task prompt vectors from similar problems are more similar.
Additionally, we evaluate the similarity of different task prompt vectors across different tasks. Figure 3 shows the cosine similarity between task prompt vectors for different tasks. We can see that certain pairs of tasks are more similar than others, what can be shared properties of these tasks, such as the same number of classes, same labels, or solving the same problem. Problem similarity can be seen in DBPedia–TREC and MNLI–QNLI task prompt vectors, and the similarity in the number of classes can be seen in the MNLI task prompt vector which tends to have higher cosine similarity with task prompt vectors for tasks with more classes (e.g., DBPedia, TREC Coarse).
4.3 Combination of Task Prompt Vectors via Arithmetic Addition for Multi-Task Transfer

### v2

@@ Abstract
Prompt tuning is an efficient solution for training large language models (LLMs). However, current soft-prompt-based methods often sacrifice multi-task modularity, requiring the training process to be fully or partially repeated for each newly added task. While recent work on task vectors applied arithmetic operations on full model weights to achieve the desired multi-task performance, a similar approach for soft-prompts is still missing. To this end, we introduce Task Prompt Vectors, created by element-wise difference between weights of tuned soft-prompts and their random initialization. Experimental results on 12 NLU datasets show that task prompt vectors can be used in low-resource settings to effectively initialize prompt tuning on similar tasks. In addition, we show that task prompt vectors are independent of the random initialization of prompt tuning on 2 different language model architectures. This allows prompt arithmetics with the pre-trained vectors from different tasks. In this way, we provide a competitive alternative to state-of-the-art baselines by arithmetic addition of task prompt vectors from multiple tasks.
@@ 1 Introduction
Figure 1: An illustration of task prompt vector and the combination via addition that we include in our work. (a) A task prompt vector is created by subtracting the soft-prompt initialization weights $\theta_{P_{pre}}$ from the soft-prompt weights after prompt tuning $\theta^{t}_{P_{ft}}$ (Section 3, eq. 3). (b) A combination via the addition of two task prompt $\tau_{P_{a}}$ and $\tau_{P_{b}}$ resulting in $\tau_{P_{new}}$ (Section 3, eq. 3). (c) Different task prompt vectors point into the same sub-space in the embedding space of PLM (Section 4.2). The circles represent different random initializations.
@@
To address the problem of the increasing number of parameters, Parameter-Efficient Fine-Tuning (PEFT) methods (Lester et al., 2021; Houlsby et al., 2019; Hu et al., 2022) were introduced, capable of solving multiple problems even with small amounts of labeled data while training only a fraction of the model parameters (e.g., for RoBERTa base (Liu et al., 2019), prompt tuning Lester et al. (2021) is training only 0.5% parameters, and LoRA (Hu et al., 2022) is training only 0.7% of parameters (Xu et al., 2023)). The key concept that makes such methods effective is their task modularity.
@@
Some of the recent PEFT (Xu et al., 2023; Lester et al., 2021; Asai et al., 2022) methods focus on fine-tuning soft-prompts. Soft-prompts are trainable (parametrized) weights that are prepended to the input embeddings while training the model. Prompt tuning is one such efficient solution for soft-prompt-based tuning of large language models (LLMs).
@@
The most of the soft-prompt-based methods lack sufficient multi-task modularity, requiring the training process to be fully or partially repeated for each newly added task (Vu et al., 2022; Wang et al., 2023). Other methods while keeping their relatively high modularity usually lack robustness and their performance depends on the quality and the number of pre-trained soft-prompts (Asai et al., 2022). Moreover, creating a soft-prompt for multiple tasks may often reduce the overall multi-task performance and require further fine-tuning. Building upon the findings from task vector arithmetics Ilharco et al. (2022), we utilize the efficiency and modularity of prompt tuning Lester et al. (2021) to create Task Prompt Vectors. We thoroughly investigate the properties of task prompt vectors and demonstrate their functionality in combining pairs of task prompt vectors while evaluating their in-distribution performance and out-of-distribution performance in full and limited data scenarios.
@@
Our main contributions and findings are  [^1] To support the replicability of our work, we provide a repository where we store all of our implementation and results: https://github.com/kinit-sk/task-prompt-vectors:
@@ 2 Related Work
To the best of our knowledge, there is no research on task vectors in the context of soft-prompt-based fine-tuning. In this work, we address this drawback by building on the existing knowledge on prompt tuning and task vectors.
3 Task Prompt Vectors
Background.
Prompt tuning, as introduced in Lester et al. (2021), casts tasks as text generation, modeling a probability $Pr(Y|X)$, where $X$ is a sequence of input tokens and $Y$ is a sequence of output tokens representing the class label. The classification $Pr_{\theta}(Y|X)$ is then parametrized by the model weights $\theta$. Prompting adds extra information to the classification process by prepending a series of tokens (prompt) $P$ to the input $X$, such that the model maximizes the probability of getting current $Y$ in $Pr_{\theta}(Y|[P;X])$, while keeping the parameters $\theta$ frozen. Prompt tuning adds another parameter $\theta_{P}$ to the equation, which parametrizes the prompt. During the training, only $\theta_{P}$ is updated as the $\mathcal{L}_{PT}=-\sum_{i}logPr_{\theta,\theta_{P}}(Y_{i}|[P;X_{i}])$ (1) function is optimized.
As a method of adapting model weights without training, task vectors (Ilharco et al., 2022) compute the difference between pre-trained weights and fine-tuned weights on a specific task. The task vector is simply defined by the element-wise difference between the pre-trained weights of the whole model and the weights after fine-tuning. Task vectors can be then applied to any model weights $\theta$ of the same dimensionality (architecture) by element-wise addition. The representation of task vectors in the weight space of the model has the same properties as standard vectors, therefore it is possible to include them in arithmetic expressions like addition, negation, or combinations via the addition of two or more vectors together. We further build on the findings from Ilharco et al. (2022) and Lester et al. (2021) in the following parts of this section.
@@ 3 Task Prompt Vectors
Let $T_{1},...,T_{t}$ be a set of source tasks and $\theta_{P_{1}},...,\theta_{P_{i}}$be a set of random soft-prompt weights initializations. Intuitively, the random soft-prompt weights initializations are random points in the embedding space of the PLM. We then move each of these points (via prompt tuning) into a task sub-space where the optimization function from the equation 3 returns the (sufficiently) minimal value and we repeat this for each task $t\in T$. These points are further denoted as task prompts – soft-prompts fine-tuned by prompt tuning to a set of downstream tasks. The straight trajectory from the initial point to the task prompt is our task prompt vector (see Figure 1 part a)).
Let $\theta_{P_{pre}}\in\mathbb{R}^{d}$ be the weights of the soft-prompt randomly initialized from the embedding vocabulary of a PLM, and $\theta^{t}_{P_{ft}}\in\mathbb{R}^{d}$ be the weights of the soft-prompt $P$ fine-tuned on a specific task $t$, using the standard prompt tuning formula. We formulate the task prompt vector $\tau_{P_{t}}$ for soft-prompt $P$ and task as an element-wise difference:$\tau_{P_{t}}=\theta^{t}_{P_{ft}}-\theta_{P_{pre}}$ (2).
Applying a task prompt vector to the soft-prompt weights of the same size would follow: $\theta_{P_{new}}=\theta_{P}+\lambda\theta^{t}_{P_{ft}}$ (3). Where the rescaling term $\lambda$ is a number from the same interval $0<\lambda\leq 1$ and when $\lambda=1$, then $\theta_{P_{new}}=\theta^{t}_{P_{ft}}=\theta_{P}$.
Vector arithmetics with task prompt vectors.
The task prompt vectors for different tasks can be combined by simple vector addition, combining knowledge from different tasks. When we experiment with combinations, we refer to the arithmetic addition of two task prompt vectors (see Figure 1 part b)): $\tau_{P_{new}}=\tau_{P_{a}}+\tau_{P_{b}}$ (4).
@@ 4.1 Experimental Setup and Implementation Details
We investigate the properties of task prompt vectors using the representative foundation T5-base (Raffel et al., 2020) model for all of our experiments and representative open autoregressive LLaMa-3.1-8B-Instruct (Dubey et al., 2024) model for origin dependency experiments in Section 4.2. Our investigation covers 3 types of classification problems and 12 corresponding datasets, namely natural language inferenceMNLI (Williams et al., 2018), QNLI (Wang et al., 2018), SciTail (Khot et al., 2018), SNLI (Bowman et al., 2015); topic classificationDBPedia (Auer et al., 2007), TREC (Li and Roth, 2002; Hovy et al., 2001), AG News, Yahoo Answers (Zhang et al., 2015); and sentiment classificationSST2 (Socher et al., 2013), Yelp Polarity, SST5, IMDB (Maas et al., 2011).
For all experimental results, we report F1 macro, if not specified otherwise. The cosine similarity between vectors (task prompts or task prompt vectors) is measured using the flattened weights of each vector (which has a size of $100\times 768$ parameters, resulting in a 76800-dimensional vector). We average our zero- and few-shot results across 3 different runs (i.e., different random initializations of soft-prompts) for ATTEMPT and multi-task SPoT baselines (mostly to save more computational resources) and across 10 different runs for all other experiments. To determine the statistical significance of our results we perform a two-sample Student’s t-test (Student, 1908) with Bonferroni correction Dunn (1959) between the best result and the second best result. If the population sizes differ (e.g. 10 and 3 runs) we use Welch’s t-test (Welch, 1947).
For the few-shot experiments (simulating limited labeled data scenarios), we randomly sub-sample from the data for the respective number of shots while keeping the class distribution. We consider shot and sample to be equivalent (i.e., for a 5-shot setting, we choose 5 samples overall, not 5 samples per class). A detailed description of our experimental setup can be found in Appendix A.
@@ 4.2 Investigating Task Prompt Vectors Properties
model dataset QNLI MNLI TREC DBpedia SST2 Yelp avg
T5 Original init $93.3_{0}$ $85.4_{0.1}$ $95.5_{1.7}$ $99.1_{0}$ $93.8_{0.3}$ $97.2_{0}$ $93.8_{0.4}$
Mixed init ${93.2_{0.1}}^{*}$ ${85.3_{0.2}}$ ${26.5_{18.2}}^{*}$ ${99_{0.1}}^{*}$ ${93.2_{0.6}}$ ${97.1_{0.1}}^{*}$ ${82.4_{3.2}}^{*}$
LLaMa Original init $92.0_{0}$ $89.7_{0.2}$ $95.8_{0.3}$ $99.2_{0}$ $95.9_{0.4}$ $98.6_{0.1}$ $95.2_{0.2}$
Mixed init $92.0_{0.1}$ $89.7_{0.2}$ $96.0_{0.3}$ $99.2_{0}$ $96.0_{0.5}$ $98.6_{0.1}$ $95.3_{0.2}$
Table 1: Comparison of exact match results across 10 random soft-prompt initializations for T5-base model and 3 initializations for LLaMa-3.1-8B model. The subscript represents the standard deviation from the average. The first row (Original init) represents the results of prompt tuning. The second row (Mixed init) represents the results of moving a specific initialization in the direction of a task prompt vector created from different (mixed) initializations.
RQ1: How universally can we apply task prompt vectors to a) different prompt initializations and b) different tasks?
@@
To evaluate these properties, we train a set of soft-prompts on specified source tasks for inference classification (MNLI, QNLI), topic classification (DBPedia, TREC), and sentiment classification (SST2, Yelp Polarity), resulting in a set of six soft-prompts that were trained from a single random initialization. We sample 10 random initializations for T5-base model from which we create the task prompt vectors as described in equation 3. For LLaMa-3.1-8B-instruct we sample only 3 random initialization to preserve computational resources. We aggregate by averaging our results across random initializations in Table 1 and Figures 2, 3. We start with the evaluation of whether the task prompt vectors are independent of the random initialization and continue with the experiments to confirm whether the trained task prompts from prompt tuning end up in the same task sub-space of the PLM embedding space. This helps us determine whether the task prompt vectors point in the same space, similar to Figure 1 part c).
@@
We conduct experiments to evaluate the performance of applying task prompt vectors to different (mixed) random initializations. For each task and each random initialization, we apply the task prompt vector (according to the equation 3) to all of the other random initializations and evaluate performance for each task prompt vector-initialization pair on the test set of the particular dataset. The aggregated results in "Mixed init" rows in Table 1 differ only slightly in most observed tasks for both of the observed models, compared to the results of prompt tuning in the "Original init" rows. This indicates that task prompt vectors perform well irrespective of their initialization. The only exception is in the TREC task, where the performance decreases drastically only for the T5-base model. We suspect that this may be caused by the task being harder for the T5-base model to learn, which also confirms the higher standard deviation from the mean of prompt tuning performance. We can also see that in the case of the LLaMa-3.1-8B-Instruct model, there is no statistically significant difference between using the original initialization or different task prompt vector initialization and for the TREC and SST2 the average results slightly increased.
Figure 2: Comparison of average cosine similarities of task prompts fine-tuned on different tasks for T5-base model. The average is calculated across all combinations of 10 random initializations (i.e., row QNLI column MNLI was calculated as the average of all cosine similarities between MNLI and QNLI task prompts for all random initialization combinations omitting the combinations where cosine similarity is equal to 1). The diagonal represents the cosine similarities of the same tasks and it represents the maximum value of cosine similarity across different random initializations.
Figure 3: Comparison of average cosine similarities of task prompt vectors. The averages are calculated equivalently to Figure 2 but with task prompt vectors created from different task prompts.
@@
Figures 2 and 3 show the comparison of cosine similarities between task prompts and task prompt vectors from different tasks averaged over all random initialization combinations. We can see from the low cosine similarities in both tables, that the task prompts and task prompt vectors do not end up in the same direction when initialized from different points in the embedding space. The highest cosine similarities on the diagonal represent the highest cosine similarity, which serves as a baseline for comparison with the cross-task cosine similarities. We can see in Table 1 row 1, that the downstream performance of prompt tuning on the source tasks across 10 different random initializations has a low standard deviation from the average. This means that the task prompts after prompt tunings end up in a subspace with sufficient task performance, without necessarily pointing to the same spot in the task subspace. Cosine similarities that we have used to create the aggregated figures can be seen in the Appendix B in Figures 6 and 8. We have also evaluated cosine similarities of task prompt vectors created from 3 different random initializations for LLaMa-3.1-8B-Instruct in Appendix B in Figures 11 and 12.hTask prompt vectors from similar problems are more similar. Additionally, we evaluate the similarity of different task prompt vectors across different tasks. Figure 3 shows the cosine similarity between task prompt vectors for different tasks. We can see that certain pairs of tasks are more similar than others, what can be shared properties of these tasks, such as the same number of classes, same labels, or solving the same problem. Problem similarity can be seen in DBPedia–TREC and MNLI–QNLI task prompt vectors, and the similarity in the number of classes can be seen in the MNLI task prompt vector which tends to have higher cosine similarity with task prompt vectors for tasks with more classes (e.g., DBPedia, TREC).
4.3 Combination of Task Prompt Vectors via Addition for Multi-Task Transfer